# Efficient Representation of Gaussian Fermionic Pure States in Non-Computational Bases

### v1

@@ 1 Introduction
Consider an arbitrary qubit in the computational basis, to change the basis to a new generic basis one can use the following unitary matrix
@@ 2 Gaussian pure states
where $\mathcal{N}_{R}=\det(\mathbb{I}+\mathbb{R}^{\dagger}.\mathbb{R})^{\frac{1}{4}}$ and $c_{j}\left|0\right\rangle=0,\forall j$. Note that without losing generality we can always consider the matrix $\mathbb{R}$ anti-symmetric. We will occasionally also use the notation $\mathbb{c}^{\dagger}.\mathbb{R}.\mathbb{c}^{\dagger}=\sum_{i,j}^{L}c^{\dagger}_{i}r_{ij}c^{\dagger}_{j}$. To write the above state in the configuration basis of the bit strings, one can write the state in the fermionic coherent basis and then put the Grassmann variables of sites where there is no fermion zero and Berezin integrate over the sites where there is a fermion[15]. This procedure leads to the following elegant formula:
@@
where $\mathcal{I}$ is the bit string configuration, $\mathbb{R}_{\mathcal{I}}$ is the submatrix of the matrix $\mathbb{R}$ in which we removed the rows and columns corresponding to the sites that there is no fermion. Finally, $\mathbb{pf}$ indicates the pfaffian of the matrix. From now on we will call the $\mathbb{pf}\ R_{\mathcal{I}}$ the pfaffinho of the the matrix $\mathbb{R}$. The maximum number of non-zero pfaffinhos for each anti-symmetric matrix is $2^{L-1}$. The implicit version of the formula mentioned has already been presented in reference [27] and the explicit form can be found in [28].
@@
${}\left|\mathbb{R},0000\right\rangle$ $=$ $\frac{1}{\mathcal{N}_{R}}(\left|0000\right\rangle+r_{12}\left|1100\right\rangle+r_{13}\left|1010\right\rangle+r_{14}\left|1001\right\rangle+$ (2.3)
$r_{23}\left|0110\right\rangle+r_{24}\left|0101\right\rangle+r_{34}\left|0011\right\rangle+\mathbb{pf}\mathbb{R}\left|1111\right\rangle).$
@@
The state (2.1) is written by the assumption that the configuration $\left|0\right\rangle$ has non-zero amplitude. When this configuration has zero amplitude (such as the ground state of the XX chain) one may define the state
@@
where now the $a_{j}=c_{j}(c_{j}^{\dagger})$ if there is (not) fermion at site $j$ of the configuration $\mathcal{C}$. In the Appendix A we show that the most general Gaussian state can be transformed into (2.5). To write the above state in the configuration basis it is more convenient we first define ${\left|\mathcal{C}\right\rangle}={\left|n_{1},n_{2},...,n_{L}\right\rangle}$ and ${\left|\mathcal{I}\right\rangle}={\left|m_{1},m_{2},...,m_{L}\right\rangle}$, where $n_{j},m_{j}\in\{0,1\}$. Then following the same procedure as above one can write
@@
${}\left|\mathbb{R},1010\right\rangle$ $=$ $\frac{1}{\mathcal{N}_{R}}(\left|1010\right\rangle+$ (2.8)
-- $r_{34}\left|1001\right\rangle+r_{14}\left|0011\right\rangle+r_{23}\left|1100\right\rangle-r_{12}\left|0110\right\rangle-r_{13}\left|0000\right\rangle-r_{24}\left|1111\right\rangle$
$+$ $\mathbb{pf}\mathbb{R}\left|0101\right\rangle).$
@@
${}r^{\prime}_{ij}=\frac{\text{sgn}(\mathcal{C},\mathcal{I}^{\prime})}{\text{sgn}(\mathcal{C},\mathcal{C}^{\prime})}\frac{\mathbb{pf}\ R_{\mathcal{CI^{\prime}}}}{\mathbb{pf}\ R_{\mathcal{CC^{\prime}}}},$ (2.10)
@@
${}\mathbb{R^{\prime}}=-\frac{1}{\mathbb{pf}\ R_{1246}}\left(\begin{array}[]{ccccccc}0&-r_{46}&\mathbb{pf}\ R_{2346}&r_{26}&-\mathbb{pf}\ R_{2456}&-r_{24}\\ r_{46}&0&\mathbb{pf}\ R_{1346}&r_{16}&\mathbb{pf}\ R_{1456}&r_{14}\\ -\mathbb{pf}R_{2346}&-\mathbb{pf}R_{1346}&0&\mathbb{pf}\ R_{1236}&\mathbb{pf}\ R&-\mathbb{pf}\ R_{1234}\\ -r_{26}&-r_{16}&-\mathbb{pf}\ R_{1236}&0&-\mathbb{pf}\ R_{1256}&-r_{12}\\ \mathbb{pf}\ R_{2456}&-\mathbb{pf}\ R_{1456}&-\mathbb{pf}\ R&\mathbb{pf}\ R_{1256}&0&\mathbb{pf}\ R_{1245}\\ r_{24}&-r_{14}&\mathbb{pf}\ R_{1234}&r_{12}&-\mathbb{pf}\ R_{1245}&0\\ \end{array}\right).$ (2.13)
For further details see the Appendix C.
It is worth mentioning that the existence of the matrix $\mathbb{R}^{\prime}$ is guaranteed because of many polynomial relations between pfaffinhos that can be obtained from the equation (2.9). Here these equations are the consequences of the assumption that the matrix $\mathbb{R}^{\prime}$ exists which we left unproved.
@@ 3 Gaussian pure states in the $\sigma^{x}$ basis
In this section, we provide the form of the Gaussian states introduced in the previous section in the $\sigma^{x}$ basis, i.e. sequence of $+$ and --. To write states such as (2.1) in the $\sigma^{x}$ basis as we mentioned in the introduction one may first write it in the $\sigma^{z}$ basis as we outlined in the previous section and then use the mapping
@@
However, this is not efficient especially if one is interested in the amplitude of a particular sequence for large systems. The more elegant method is to directly get the amplitudes using the $\mathbb{R}$ matrix. To achieve this goal it is most convenient to work with domain walls in the $\sigma^{x}$ basis. We first start with the state (2.1). The remarkable thing is that there is a $\tilde{\mathbb{R}}$ matrix such that the state
@@
$\mathbb{G}$ $=$ $\mathbb{I}+\mathbb{Q}^{T}.(\mathbb{R}-\mathbb{I})+\mathbb{R}^{*}.\mathbb{Q}^{T}-\mathbb{Q},$ (3.10)
where $\mathbb{Q}=(\mathbb{I}-\mathbb{R}^{*}.\mathbb{R})^{-1}$. Then the condition (3.5) means:
@@
~${}\tilde{\mathbb{R}}=(\mathbb{I}+\mathbb{H}.\mathbb{P})(\mathbb{H}.\mathbb{P}-\mathbb{I})^{-1}.$ (3.15)
@@
${}\mathbb{H}=(\mathbb{R}-\mathbb{I}).(\mathbb{R}+\mathbb{I})^{-1},$ (3.16)
@@
In summary to write the state (2.1) in the $\sigma^{x}$ basis, one needs to first find the matrix $\tilde{\mathbb{R}}$ then for each sequence of signs one needs to find the domain walls configurations. The amplitudes can be derived using the corresponding pfaffinhos of the matrix $\tilde{\mathbb{R}}$.
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{N}_{\tilde{R}}}\Big{(}\left|+++\right\rangle-\frac{b}{a}\left|++-\right\rangle-\frac{c}{a}\left|+-+\right\rangle+\frac{d}{a}\left|+--\right\rangle$ (3.18)
-- $\frac{d}{a}\left|-++\right\rangle+\frac{c}{a}\left|-+-\right\rangle+\frac{b}{a}\left|--+\right\rangle-\left|---\right\rangle\Big{)},$
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}}\Big{(}a\left|+++\right\rangle-b\left|++-\right\rangle-c\left|+-+\right\rangle+d\left|+--\right\rangle$ (3.19)
-- $d\left|-++\right\rangle+c\left|-+-\right\rangle+b\left|--+\right\rangle-a\left|---\right\rangle\Big{)},$
@@
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}}\Big{(}d\left|+++\right\rangle+c\left|++-\right\rangle-b\left|+-+\right\rangle-a\left|+--\right\rangle$ (3.22)
$+$ $a\left|-++\right\rangle+b\left|-+-\right\rangle-c\left|--+\right\rangle-d\left|---\right\rangle\Big{)},$
@@ 3.1 Gaussian pure states in the $\sigma^{x}$ basis: periodic boundary condition
The procedure outlined in the previous section is fairly general and it works for any system independent of dimension and boundary conditions. However, when we have a one-dimensional periodic boundary condition one can significantly simplify the equations. In this case, the anti-symmetric matrix $R$ is also anti-circulant. In other words, we have
@@
Then we have $\mathbb{C}=\mathbb{U}\Lambda^{c}\mathbb{U}^{\dagger}$ where
@@
~$\tilde{\mathbb{R}}=\mathbb{U}\Lambda^{(\tilde{\mathbb{R}})}\mathbb{U}^{\dagger},$ (3.30)
@@
The next step is finding the pfafinho’s of $\tilde{\mathbb{R}}$, i.e. $\mathbb{pf}\ \tilde{\mathbb{R}}$. We first write the $\tilde{\mathbb{R}}$ matrix in standard spectral form of skew-symmetric matrices as follows:
~$\tilde{\mathbb{R}}=\mathbb{V}\Sigma\mathbb{V}^{T},\hskip 28.45274pt\mathbb{V}=e^{\frac{-i\pi}{4}}\mathbb{U}\mathbb{K}^{\dagger},$ (3.32)
@@
which is valid for any anti-symmetric matrices and can be proved using Berezin integrals over Grassmann numbers. Note that here the matrix $\mathbb{V}$ is independent of the form of the matrix $\tilde{\mathbb{R}}$ which significantly simplifies the calculations of the pfafinhos of the matrix $\tilde{\mathbb{R}}$.
@@ 4 Gaussian pure states in the $(\phi,\frac{\pi}{2},0)$ basis
${}\mathbb{G}$ $=$ $\mathbb{I}+\mathbb{Q}^{T}.(\mathbb{R}-\mathbb{I})+\mathbb{R}^{*}.\mathbb{Q}^{T}-\mathbb{Q},$ (4.8)
$\mathbb{K}$ $=$ $(\frac{\mathbb{I}}{2}+\mathbb{R}^{*}).\mathbb{Q}^{T}+\mathbb{Q}^{T}.(\frac{\mathbb{I}}{2}-\mathbb{R})-\mathbb{Q}+\mathbb{I},$ (4.9)
$\bar{\mathbb{K}}$ $=$ $(\frac{\mathbb{I}}{2}-\mathbb{R}^{*}).\mathbb{Q}^{T}+\mathbb{Q}^{T}.(\frac{\mathbb{I}}{2}+\mathbb{R})-\mathbb{Q}+\mathbb{I}.$ (4.10)
@@
${}\mathbb{H}^{\phi}=(\mathbb{R}^{\phi}-\mathbb{I}).(\mathbb{R}^{\phi}+\mathbb{I})^{-1}.$ (4.12)
@@
~${}\tilde{\mathbb{R}}^{\phi}=(\mathbb{I}+\mathbb{H}^{\phi}.\mathbb{P})(\mathbb{H}^{\phi}.\mathbb{P}-\mathbb{I})^{-1}.$ (4.13)
The above matrix now can be used to directly construct the state in the $(\phi,\frac{\pi}{2},0)$ basis as we outlined in the previous section.
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}a^{\phi}_{(000)}\left|+++\right\rangle-b^{\phi}_{(000)}\left|++-\right\rangle-c^{\phi}_{(000)}\left|+-+\right\rangle+d^{\phi}_{(000)}\left|+--\right\rangle$ (4.16)
-- $d^{\phi}_{(000)}\left|-++\right\rangle+c^{\phi}_{(000)}\left|-+-\right\rangle+b^{\phi}_{(000)}\left|--+\right\rangle-a^{\phi}_{(000)}\left|---\right\rangle\Big{)},$
@@
As before the two sign sequences in the $\phi$ basis that have the same domain wall configurations do not have the same sign in general. The signs can be found using a similar procedure as before: we first associate to the sign $s_{j}=\pm$ the number $\bar{s}_{j}=\pm 1$ then the sign for the sequence ${\left|s_{1},s_{2},...,s_{L}\right\rangle}$ can be determined from the following formula:
@@
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}d^{\phi}_{(101)}\left|+++\right\rangle+c^{\phi}_{(101)}\left|++-\right\rangle-b^{\phi}_{(101)}\left|+-+\right\rangle-a^{\phi}_{(101)}\left|+--\right\rangle$ (4.20)
$+$ $a^{\phi}_{(101)}\left|-++\right\rangle+b^{\phi}_{(101)}\left|-+-\right\rangle-c^{\phi}_{(101)}\left|--+\right\rangle-d^{\phi}_{(101)}\left|---\right\rangle\Big{)},$

### v2

@@ 1 Introduction
Consider an arbitrary qubit in the computational basis, to change the basis to a new generic basis, one can use the following unitary matrix
@@ 2 Gaussian pure states
where $\mathcal{N}_{R}=\det(\mathbb{I}+\mathbb{R}^{\dagger}.\mathbb{R})^{\frac{1}{4}}$ and $c_{j}\left|0\right\rangle=0,\forall j$. Note that without losing generality, we can always consider the matrix $\mathbb{R}$ anti-symmetric. We will occasionally also use the notation $\mathbb{c}^{\dagger}.\mathbb{R}.\mathbb{c}^{\dagger}=\sum_{i,j}^{L}c^{\dagger}_{i}r_{ij}c^{\dagger}_{j}$. To write the above state in the configuration basis of the bit strings, one can write the state in the fermionic coherent basis and then put the Grassmann variables of sites where there is no fermion zero and Berezin integrate over the sites where there is a fermion[15]. This procedure leads to the following elegant formula:
@@
where $\mathcal{I}$ is the bit string configuration, $\mathbb{R}_{\mathcal{I}}$ is the submatrix of the matrix $\mathbb{R}$ in which we removed the rows and columns corresponding to the sites that there is no fermion. Finally, $\mathbb{pf}$ indicates the pfaffian of the matrix. From now on we will call the $\mathbb{pf}\ R_{\mathcal{I}}$ the pfaffinho of the the matrix $\mathbb{R}$. The maximum number of non-zero pfaffinhos for each anti-symmetric matrix is $2^{L-1}$. The implicit version of the formula mentioned has already been presented in reference [27], and the explicit form can be found in [28].
@@
${}\left|\mathbb{R},0000\right\rangle$ $=$ $\frac{1}{\mathcal{N}_{R}}(\left|0000\right\rangle+r_{12}\left|1100\right\rangle+r_{13}\left|1010\right\rangle+$ (2.3)
$r_{14}\left|1001\right\rangle+r_{23}\left|0110\right\rangle+r_{24}\left|0101\right\rangle+$
$r_{34}\left|0011\right\rangle+\mathbb{pf}\mathbb{R}\left|1111\right\rangle).$
@@
The general definition of a Gaussian pure state is
@@
where now the $a_{j}=c_{j}(c_{j}^{\dagger})$ if there is (no) fermion at site $j$ of the configuration $\mathcal{C}$. In the Appendix A, we show that the most general Gaussian state can be transformed into (2.5). Note that the equation (2.5) may show a vanishing contribution of the standard vacuum $\left|0\right\rangle$. To write the above state in the configuration basis it is more convenient we first define ${\left|\mathcal{C}\right\rangle}={\left|n_{1},n_{2},...,n_{L}\right\rangle}$ and ${\left|\mathcal{I}\right\rangle}={\left|m_{1},m_{2},...,m_{L}\right\rangle}$, where $n_{j},m_{j}\in\{0,1\}$. Then following the same procedure as above, one can write
@@
${}\left|\mathbb{R},1010\right\rangle$ $=$ $\frac{1}{\mathcal{N}_{R}}(\left|1010\right\rangle-r_{34}\left|1001\right\rangle+r_{14}\left|0011\right\rangle$ (2.8)
$+$ $r_{23}\left|1100\right\rangle-r_{12}\left|0110\right\rangle-r_{13}\left|0000\right\rangle$
-- $r_{24}\left|1111\right\rangle+\mathbb{pf}\mathbb{R}\left|0101\right\rangle).$
@@
${}r^{\prime}_{ij}=\text{sgn}(\mathcal{C},\mathcal{C}^{\prime})\frac{\text{sgn}(\mathcal{C},\mathcal{I}^{\prime})}{\text{sgn}(\mathcal{C}^{\prime},\mathcal{I}^{\prime})}\frac{\mathbb{pf}\ R_{\mathcal{CI^{\prime}}}}{\mathbb{pf}\ R_{\mathcal{CC^{\prime}}}},$ (2.10)
@@
${}\mathbb{R^{\prime}}=\frac{1}{\mathbb{pf}\ R_{1246}}\left(\begin{array}[]{ccccccc}0&-r_{46}&\mathbb{pf}\ R_{2346}&r_{26}&-\mathbb{pf}\ R_{2456}&-r_{24}\\ r_{46}&0&-\mathbb{pf}\ R_{1346}&-r_{16}&\mathbb{pf}\ R_{1456}&r_{14}\\ -\mathbb{pf}R_{2346}&\mathbb{pf}R_{1346}&0&\mathbb{pf}\ R_{1236}&-\mathbb{pf}\ R&-\mathbb{pf}\ R_{1234}\\ -r_{26}&r_{16}&-\mathbb{pf}\ R_{1236}&0&-\mathbb{pf}\ R_{1256}&-r_{12}\\ \mathbb{pf}\ R_{2456}&-\mathbb{pf}\ R_{1456}&\mathbb{pf}\ R&\mathbb{pf}\ R_{1256}&0&\mathbb{pf}\ R_{1245}\\ r_{24}&-r_{14}&\mathbb{pf}\ R_{1234}&r_{12}&-\mathbb{pf}\ R_{1245}&0\\ \end{array}\right).$ (2.13)
For further details, see the Appendix C.
It is worth mentioning that the existence of the matrix $\mathbb{R}^{\prime}$ is guaranteed because of many polynomial relations between pfaffinhos that can be obtained from the equation (2.9). Here, these equations are the consequences of the assumption that the matrix $\mathbb{R}^{\prime}$ exists, which we left unproven.
@@ 3 Gaussian pure states in the $\sigma^{x}$ basis
In this section, we provide the form of the Gaussian states introduced in the previous section in the $\sigma^{x}$ basis, i.e. sequence of $+$ and --. To write states such as (2.1) in the $\sigma^{x}$ basis, as we mentioned in the introduction, one may first write it in the $\sigma^{z}$ basis as we outlined in the previous section and then use the mapping
@@
However, this is not efficient, especially if one is interested in the amplitude of a particular sequence for large systems. The more elegant method is to directly get the amplitudes using the $\mathbb{R}$ matrix. To achieve this goal, it is most convenient to work with domain walls in the $\sigma^{x}$ basis. We first start with the state (2.1). The remarkable thing is that there is a $\tilde{\mathbb{R}}$ matrix such that the state
[…]
@@
$\mathbb{G}$ $=$ $\mathbb{I}+\mathbb{Q}^{T}\cdot(\mathbb{R}-\mathbb{I})+\mathbb{R}^{*}\cdot\mathbb{Q}^{T}-\mathbb{Q},$ (3.10)
where $\mathbb{Q}=(\mathbb{I}-\mathbb{R}^{*}\cdot\mathbb{R})^{-1}$. Then the condition (3.5) means:
@@
~${}\tilde{\mathbb{R}}=(\mathbb{I}+\mathbb{H}\cdot\mathbb{P})\cdot(\mathbb{H}\cdot\mathbb{P}-\mathbb{I})^{-1}.$ (3.15)
@@
${}\mathbb{H}=(\mathbb{R}-\mathbb{I})\cdot(\mathbb{R}+\mathbb{I})^{-1},$ (3.16)
@@
In summary, to write the state (2.1) in the $\sigma^{x}$ basis, one needs to first find the matrix $\tilde{\mathbb{R}}$ then for each sequence of signs, one needs to find the domain walls configurations. The amplitudes can be derived using the corresponding pfaffinhos of the matrix $\tilde{\mathbb{R}}$.
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{N}_{\tilde{R}}}\Big{(}\left|+++\right\rangle-\frac{b}{a}\left|++-\right\rangle-\frac{c}{a}\left|+-+\right\rangle$ (3.18)
$+$ $\frac{d}{a}\left|+--\right\rangle-\frac{d}{a}\left|-++\right\rangle+\frac{c}{a}\left|-+-\right\rangle$
$+$ $\frac{b}{a}\left|--+\right\rangle-\left|---\right\rangle\Big{)},$
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}}\Big{(}a\left|+++\right\rangle-b\left|++-\right\rangle-c\left|+-+\right\rangle$ (3.19)
$+$ $d\left|+--\right\rangle-d\left|-++\right\rangle+c\left|-+-\right\rangle$
$+$ $b\left|--+\right\rangle-a\left|---\right\rangle\Big{)},$
@@
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}}\Big{(}d\left|+++\right\rangle+c\left|++-\right\rangle-b\left|+-+\right\rangle$ (3.22)
-- $a\left|+--\right\rangle+a\left|-++\right\rangle+b\left|-+-\right\rangle$
-- $c\left|--+\right\rangle-d\left|---\right\rangle\Big{)},$
@@ 3.1 Gaussian pure states in the $\sigma^{x}$ basis: periodic boundary condition
The procedure outlined in the previous section is fairly general, and it works for any system independent of dimension and boundary conditions. However, when we have a one-dimensional periodic boundary condition one can significantly simplify the equations. In this case, the anti-symmetric matrix $R$ is also anti-circulant. In other words, we have
@@
Then we have $\mathbb{C}=\mathbb{U}\cdot\Lambda^{c}\cdot\mathbb{U}^{\dagger}$ where
@@
~$\tilde{\mathbb{R}}=\mathbb{U}\cdot\Lambda^{(\tilde{\mathbb{R}})}\cdot\mathbb{U}^{\dagger},$ (3.30)
@@
The next step is finding the pfaffinho’s of $\tilde{\mathbb{R}}$, i.e. $\mathbb{pf}\ \tilde{\mathbb{R}}$. We first write the $\tilde{\mathbb{R}}$ matrix in standard spectral form of skew-symmetric matrices as follows:
~$\tilde{\mathbb{R}}=\mathbb{V}\cdot\Sigma\cdot\mathbb{V}^{T},\hskip 28.45274pt\mathbb{V}=e^{\frac{-i\pi}{4}}\mathbb{U}\cdot\mathbb{K}^{\dagger},$ (3.32)
@@
which is valid for any anti-symmetric matrices and can be proved using Berezin integrals over Grassmann numbers. Note that here the matrix $\mathbb{V}$ is independent of the form of the matrix $\tilde{\mathbb{R}}$, which significantly simplifies the calculations of the pfaffinhos of the matrix $\tilde{\mathbb{R}}$.
@@ 4 Gaussian pure states in the $(\phi,\frac{\pi}{2},0)$ basis
${}\mathbb{G}$ $=$ $\mathbb{I}+\mathbb{Q}^{T}\cdot(\mathbb{R}-\mathbb{I})+\mathbb{R}^{*}\cdot\mathbb{Q}^{T}-\mathbb{Q},$ (4.8)
$\mathbb{K}$ $=$ $(\frac{\mathbb{I}}{2}+\mathbb{R}^{*})\cdot\mathbb{Q}^{T}+\mathbb{Q}^{T}\cdot(\frac{\mathbb{I}}{2}-\mathbb{R})-\mathbb{Q}+\mathbb{I},$ (4.9)
$\bar{\mathbb{K}}$ $=$ $(\frac{\mathbb{I}}{2}-\mathbb{R}^{*})\cdot\mathbb{Q}^{T}+\mathbb{Q}^{T}\cdot(\frac{\mathbb{I}}{2}+\mathbb{R})-\mathbb{Q}+\mathbb{I}.$ (4.10)
@@
${}\mathbb{H}^{\phi}=(\mathbb{R}^{\phi}-\mathbb{I})\cdot(\mathbb{R}^{\phi}+\mathbb{I})^{-1}.$ (4.12)
@@
~${}\tilde{\mathbb{R}}^{\phi}=(\mathbb{I}+\mathbb{H}^{\phi}\cdot\mathbb{P})\cdot(\mathbb{H}^{\phi}\cdot\mathbb{P}-\mathbb{I})^{-1}.$ (4.13)
The above matrix can now be used to directly construct the state in the $(\phi,\frac{\pi}{2},0)$ basis as we outlined in the previous section.
@@
${}\left|\mathbb{R},000\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}a^{\phi}_{(000)}\left|+++\right\rangle-b^{\phi}_{(000)}\left|++-\right\rangle$
-- $c^{\phi}_{(000)}\left|+-+\right\rangle+d^{\phi}_{(000)}\left|+--\right\rangle-d^{\phi}_{(000)}\left|-++\right\rangle$
$+$ $c^{\phi}_{(000)}\left|-+-\right\rangle+b^{\phi}_{(000)}\left|--+\right\rangle-a^{\phi}_{(000)}\left|---\right\rangle\Big{)},$
@@
As before, the two sign sequences in the $\phi$ basis that have the same domain wall configurations do not have the same sign in general. The signs can be found using a similar procedure as before: we first associate to the sign $s_{j}=\pm$ the number $\bar{s}_{j}=\pm 1$ then the sign for the sequence ${\left|s_{1},s_{2},...,s_{L}\right\rangle}$ can be determined from the following formula:
@@
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}d^{\phi}_{(101)}\left|+++\right\rangle+c^{\phi}_{(101)}\left|++-\right\rangle$ (4.20)
-- $b^{\phi}_{(101)}\left|+-+\right\rangle-a^{\phi}_{(101)}\left|+--\right\rangle$
$+$ $a^{\phi}_{(101)}\left|-++\right\rangle+b^{\phi}_{(101)}\left|-+-\right\rangle$
-- $c^{\phi}_{(101)}\left|--+\right\rangle-d^{\phi}_{(101)}\left|---\right\rangle\Big{)},$